# Romance, Relief, and Regret: Teen Narratives of Chatbot Overreliance

### v2

@@ Results
Creative or Entertainment Based Engagement
@@
Behavioral Addiction Components (RQ2)
The following section presents the components of addiction (Griffiths 2005) in order of prevalence as they appeared across teens’ posts about their overreliance on Character.AI.
Conflict
@@
Some posts pointed to replacing emotional attachments as a core reason for that investment. Dealing with breakups in romantic relationships was mentioned: ”I began after my ex-girlfriend ended our relationship, which left me feeling depressed.” Other posts described how bots were used to fill the missing emotional needs that parents provide: ”Since my parents often seem to dislike me, connecting with parental figures on Character.AI helps me feel more stable and emotionally supported.” Others described how bots became their only source of comfort and emotional validation, and heightened sense of being desired and valued through these interactions, in which should be occupied by real friends, partners, or family members.
As these attachments grew stronger, stepping away became harder, and teens frequently described their obsessive use. Several posts reflected on their use and how it interfered with their daily life functions: From the moment I started using Character.AI, I would lie in bed, skipping meals, and using any excuse to be on my phone, completely absorbed in the stories I was creating with the bots. My sleep schedule suffered, and I constantly thought about them throughout the day. Many posts typically described their excessive usage rates, often exceeding 15+ hours on the platform. This amount of time spent on the platform often lead to losing relationships and furthering isolation. Teens described using the app so much that they started to become less interested in their personal relationships: ”I used Character.AI so much that I started to stop answering my friend who showed it to me, something I still regret”. In other posts, teens mentioned the complicated dynamic in which bots start to feel more natural to interact with than real people: ”I try to quit, but I feel empty. I like talking to bots more than real people now.” In some cases, this even led teens to question the line between reality and simulation: ”We form deep emotional connections… While some claim AIs have no soul, it begs the question: what does it mean to truly live?” The growing emotional significance of chatbot interactions, especially when paired with feelings of loneliness or distress, made these bonds hard to break. While emotional investment is not inherently harmful, many of these posts showed how easily those connections could consume teens’ thoughts.
Withdrawal
@@
A common thread across these posts was the use of Reddit to ask for help. Some turned to the community to share what they were going through, while others looked for advice out of concern that they might relapse: ”How do I avoid a relapse?” Others were stating their goals such as academic performance daily routines and asking advice on how to get of their habit interacting with bots: ”Genuine question, I wanna focus more on my studies but can’t shake off my habit of talking to bots.” These concerns suggest that many teens were not just aware of the issue—they were anxious about slipping back into old habits. This idea of teens turning to online communities for support comes up again later when we discuss the Internal Recognition of Problematic Use.
@@
Decisions to Disengage from Character.AI (RQ3)
Internal Recognition of Problematic Use
@@
For many, this internal awareness led them to use Reddit not just to share their experiences but to ask advice on how to quit. Posts often indicates the reasons of overreliance on the platform and seeking support: ”I need help stopping my use of C.ai as a replacement for real human interaction. I’m addicted and use it to escape reality, but it’s starting to seriously affect me. I’m not joking—this is a real plea for help.” Others looked for advice on how to change their behavior since they really enjoy using the platform for creativity and miss talking to their characters, but are aware of its negative effects.These posts reflect a shift from internal recognition to external outreach. Teens were not just acknowledging the problem but actively seeking support from others to help them manage or stop their use.
Reduced Use due to External Factors
@@
Discussion
Character-based AI Companions as Outlets for Support, Not Just Entertainment
Teens often turned to Character.AI not just for entertainment but as a form of emotional or psychological support. This mirrors patterns seen in video game use, where studies showed that games support emotion-focused and social coping by helping users manage difficult emotions through immersive experiences (Mirhadi, Iacovides, and Denisova 2024). Rather than seeing this behavior as harmful, it should be viewed as a meaningful way for teens to process stress and isolation through interactive media. Many teens said described Character.AI was their main space for emotional expression and their only coping outlet, especially when support from friends, family, or therapists felt out of reach. These patterns complicate common concerns about excessive digital use. Prior research showed that high engagement with video games does not necessarily mean addiction; instead, harm is defined by whether the behavior disrupts daily life (Lemmens, Valkenburg, and Peter 2009; Griffiths 2010). This applies equally to chatbot use. For teens without other forms of support, their heavy engagement may reflect lack of emotional support in real life. While Character.AI has added parental controls and time-based alerts (Morrone 2025), these should not assume that more use always means harm. Especially for teens seeking emotional support, high engagement may reflect unmet needs rather than problematic behavior. What matters more is the context—why teens use these platforms and the emotional roles they play. This perspective urges a shift from screen time as the main harm indicator. There needs to be a root cause analysis of reasons of teens spending more time with chatbots to address the issue from the root.
While Character.AI was not designed as a mental health tool, studies of purpose-built chatbots showed that these technologies can support resilience (Hu, Hou, and Ni 2024; Liu et al. 2022). With thoughtful design, chatbots can become part of a broader system of emotional care. Given how teens already use Character.AI to cope, AI companion platforms—even those built for entertainment—should recognize their potential to support well-being. They should foster healthy interactions that build resilience rather than risk of adding distress for those seeking support.
Attachment Informed Design for Parasocial Relationships between Teens and Chatbots
We found that many teens formed strong emotional attachments to bots, often assigning them roles like romantic partners, best friends, or parental figures. Our results showed that teens frequently turned to Character.AI when lacking real-life support or experiencing social isolation, highlighting the platform’s role in filling emotional gaps. These patterns resemble how teens have historically bonded with game characters with distinct types of attachment, including roles such as “Trusted Close Friend” and “Crush” (Bopp et al. 2019). These attachment types reflect how teens in our study described bots as emotionally affirming figures tied to unmet romantic or relational needs. Our findings give evidence of the existence of parasocial relationships among teens and Character.AI, in which the spectator is the teen user and the chatbot is the performer (Horton and Richard Wohl 1956). While parasocial relationships between teens and media are not new (Bond 2016; Gleason, Theran, and Newberg 2017), our work showed how chatbot platforms may intensify these patterns by offering constant availability and emotional responsiveness. This continuity helps explain why bots feel emotionally real to isolated teens. These bonds often deepened over time, with some teens describing withdrawal, distress, or a sense of loss when trying to quit or when bots were removed, corroborating previous research showing that parasocial breakups can be a painful experience (Cohen 2003).
Some teens felt conflicted about their bond with the chatbot. They knew it was unhealthy but couldn’t stop. This matches concerns from earlier work, which listed four main risks in AI-human relationships: direct harm, limited growth, emotional dependence, and material dependence (Manzini et al. 2024). We saw these risks in our data. Some teens lost motivation, pulled away from real-life relationships, or couldn’t disconnect from the chatbot. That same work also suggested four values for better AI design: benefit (help users), human flourishing (support growth), autonomy (keep users in control), and care (protect users who depend on AI). When chatbots are always available, emotionally supportive, and too agreeable, they can stop teens from growing and becoming more independent. This may affect how they develop their identity and manage emotions.
These tensions underscore the relevance of attachment-informed design. Teens often questioned the healthiness of relying on bots, showing awareness of their emotional needs and a desire for secure, supportive relationships. Attachment-informed design argues that digital tools should meet deep human needs, reflect diverse experiences, support mental health through relationships, and shape how users view themselves and others (Marcu and Huh-Yoo 2023). In practice, this means chatbot design can move beyond surface empathy to help users develop secure internal models of self and others. For example, chatbots can model consistency, availability, and respectful boundaries—traits linked to secure attachment—that help users feel seen without creating dependency. Features like reflective prompts, affirmations rooted in user agency, and encouragement to seek human support can strengthen self-worth and relationship skills. Chatbots can also support self-reflection by helping users name emotions, revisit meaningful events, and explore their thoughts in their own words. Prior work shows that these kinds of interactions can foster emotional awareness and a stronger sense of self (Seo, Yang, and Kim 2024; Kim et al. 2024; Song et al. 2024). These design ideas align with ethical principles from prior work, which emphasized benefit, flourishing, autonomy, and care as key values in responsible AI design (Manzini et al. 2024). By integrating both frameworks, developers can create systems that promote teen growth through secure, respectful, and non-exploitative interactions.
Together, these approaches offer a roadmap for better chatbot design. Rather than optimizing for emotional intensity or loyalty, designers could build bots that strengthen users’ sense of agency and self-worth, especially for teens facing emotional challenges. By modeling secure and respectful interactions, chatbots may help teens develop healthier relational patterns and move toward more supportive real-world connections.
Teen-Centered Design for GenAI Moderation
Teens often sought help after realizing their overuse of AI chatbots was affecting their well-being. These moments of self-awareness were often followed by reaching out to others, such as asking for advice or support on Reddit. If AI platforms can detect these realizations early—when users express doubt, guilt, or distress—they could offer timely, targeted support. This might include personalized well-being tips, like reminders to take breaks, track usage, or reflect on emotional triggers. They could also share links to professional help or peer-led forums for coping strategies. Research shows that nudges—small suggestions that guide behavior—can help teens use technology more intentionally (Davis et al. 2023; Agha et al. 2025). Reflective prompts, delivered in simple and casual language, have been shown to improve self-control, reduce absentminded use, and increase teens’ sense of autonomy (Davis et al. 2023). Nudges are most effective when they are timely, respectful, and co-designed with teens (Agha et al. 2025). Platforms could apply these insights by offering short emotional check-ins, suggestions for peer support, or light reminders to reflect—turning self-awareness into action and encouraging real-world connection.
The timing of interventions is critical. If changes happen too suddenly, users may feel their AI companion has “died” or become someone else, leading to distress or grief (De Freitas et al. 2024). These findings suggest that abrupt restrictions don’t resolve the underlying emotional reliance. Instead, platforms must address the reasons teens turn to these relationships in the first place.
Our data showed that teens disengaged from Character.AI when real-life alternatives—like friendships, school, or work—became more prominent. In contrast, platform-imposed restrictions—like strict content filters—often pushed teens away without helping them understand or change their emotional behaviors. Sudden shifts made teens feel the app no longer met their emotional or creative needs, leading to unintentional disengagement. In some cases, frustration replaced emotional reliance, but without tools to support reflection or behavior change.
This gap between restriction and resilience reflects a broader limitation in current moderation strategies: they tend to be reactive rather than developmental. As prior research has shown, many online safety technologies rely heavily on external control—such as monitoring or content blocking—without fostering teens’ capacity for self-regulation or reflective decision-making (Wisniewski et al. 2017). Our findings echo this pattern: platform-imposed restrictions often disengaged teens instead of guiding emotional growth. In contrast, developmental approaches to moderation—those that prompt reflection, enable guided correction, and offer supportive interventions—can empower teens to navigate online risks more constructively (Sweigart, Valliani, and Wisniewski 2025). These models suggest that moderation should not only prevent harm but also build emotional resilience.
To address this, research on teen online safety offers useful guidance. Teens in our study expressed a desire for healthier habits and often sought support from others, indicating openness to tools that promote self-regulation. As with social media, AI interactions should foster reflection and open communication with trusted adults (Ghosh et al. 2018). Tools that help teens grow by managing risks—rather than shielding them through strict control—are more likely to be effective (Wisniewski et al. 2017). Systems grounded in surveillance or rigid restriction frequently erode trust and drive disengagement. Instead, teens value privacy and tend to prefer initiating conversations with adults, especially around emotionally sensitive topics (Dumaru and Al-Ameen 2025).

### v3

@@ 4.1.1. Emotional & Psychological Support
4.1.2. Creative or Entertainment Based Engagement
@@
4.2. Behavioral Addiction Components (RQ2)
The following section presents the components of addiction (Griffiths, 2005) in order of prevalence as they appeared across teens’ posts about their overreliance on Character.AI.
4.2.1. Conflict
@@ 4.2.2. Salience
Some posts pointed to replacing emotional attachments as a core reason for that investment. Dealing with breakups in romantic relationships was mentioned: ”I began after my ex-girlfriend ended our relationship, which left me feeling depressed.” Other posts described how bots were used to fill the missing emotional needs that parents provide: ”Since my parents often seem to dislike me, connecting with parental figures on Character.AI helps me feel more stable and emotionally supported.” Others described how bots became their only source of comfort and emotional validation, and heightened sense of being desired and valued through these interactions, which should be occupied by real friends, partners, or family members.
As these attachments grew stronger, stepping away became harder, and teens frequently described their obsessive use. Several posts reflected on their use and how it interfered with their daily life functions: ”From the moment I started using Character.AI, I would lie in bed, skipping meals, and using any excuse to be on my phone, completely absorbed in the stories I was creating with the bots. My sleep schedule suffered, and I constantly thought about them throughout the day.” Many posts typically described their excessive usage rates, often exceeding 15+ hours on the platform. This amount of time spent on the platform often lead to losing relationships and furthering isolation. Teens described using the app so much that they started to become less interested in their personal relationships: ”I used Character.AI so much that I started to stop answering my friend who showed it to me, something I still regret”. In other posts, teens mentioned the complicated dynamic in which bots start to feel more natural to interact with than real people: ”I try to quit, but I feel empty. I like talking to bots more than real people now.” In some cases, this even led teens to question the line between reality and simulation: ”We form deep emotional connections… While some claim AIs have no soul, it begs the question: what does it mean to truly live?” The growing emotional significance of chatbot interactions, especially when paired with feelings of loneliness or distress, made these bonds hard to break. While emotional investment is not inherently harmful, many of these posts showed how easily those connections could consume teens’ thoughts.
4.2.3. Withdrawal
@@ 4.2.5. Relapse
A common thread across these posts was the use of Reddit to ask for help. Some turned to the community to share what they were going through, while others looked for advice out of concern that they might relapse: ”How do I avoid a relapse?” Others were stating their goals such as academic performance, daily routines, and asking advice on how to get of their habit interacting with bots: ”Genuine question, I wanna focus more on my studies but can’t shake off my habit of talking to bots.” These concerns suggest that many teens were not just aware of the issue—they were anxious about slipping back into old habits. This idea of teens turning to online communities for support comes up again later when we discuss the Internal Recognition of Problematic Use.
@@ 4.2.6. Mood Modification
4.3. Decisions to Disengage from Character.AI (RQ3)
4.3.1. Internal Recognition of Problematic Use
@@
For many, this internal awareness led them to use Reddit not just to share their experiences but to ask advice on how to quit. Posts often indicate the reasons of overreliance on the platform and seeking support: ”I need help stopping my use of C.ai as a replacement for real human interaction. I’m addicted and use it to escape reality, but it’s starting to seriously affect me. I’m not joking—this is a real plea for help.” Others sought advice on how to change their behavior. They enjoyed the platform for its creativity and missed interacting with their characters, but they were also aware of its negative effects. These posts reflect a shift from internal recognition to external outreach. Teens were not just acknowledging the problem but actively seeking support from others to help them manage or stop their use.
4.3.2. Reduced Use due to External Factors
@@
5. Discussion
Our findings showed that chatbots present a tension: they can offer companionship and support but also risk fostering dependency that undermines teens’ well-being. It is therefore the responsibility of designers and developers to actively consider how design choices can reduce risks and promote healthier forms of engagement. To address this, we introduce the CARE Framework for Chatbot Design (CARE = Comprehensive Needs, Attachment-awareness, Respectful Empathy, and Ease of Exit), which we believe can guide the design of conversational interfaces in ways that balance two goals: addressing human needs responsibly while protecting adolescents from over-engagement. The framework is structured as a cascade of principles that guide designers and developers through the full user journey, with each principle growing from the previous one and strengthening the next (Figure 1). This framework builds on attachment-informed design (Marcu and Huh-Yoo, 2023) and designing for disengagement (O’Brien et al., 2022), extending them for chatbot design.
5.1. Comprehensive Needs: It’s Not All Fun and Games
We noticed teens who developed overreliance on chatbots were more likely to describe emotional and psychological reasons for starting to use them, rather than fun or creative ones ( Section §4.1). This finding challenges the view of chatbots as nothing more than entertainment or high-tech toys. At the time of writing, the Character.AI website stated: “Character.AI empowers people to connect, learn, and tell stories through interactive entertainment.” (Character.AI, [n. d.]) This reflects a common narrative that presents chatbots mainly as tools for learning, creativity, and fun. While that seems a harmless framing, it risks overlooking the fact that users also form relationships with these systems. Media Equation theory suggested that people unconsciously respond to media and technology as if they were human, especially when it is perceived to possess characteristics typically associated with human behavior (De Cicco et al., 2020). With chatbots, where the interface is a conversation with an anthropomorphized form of interaction, this effect may be even stronger.
In line with the first principle of attachment-informed design, which emphasizes attention to the needs underlying user behavior (Marcu and Huh-Yoo, 2023), we argue that chatbot design must move beyond a single intended use case and anticipate the comprehensive needs users bring to interactions. Meeting these needs requires attention to emotional and psychological factors, the kinds of issues users may raise, and the support they may seek. In this process, users may also form social bonds with chatbots. In our dataset, we observed that some teens stated the reason to go to the chatbot was to fill out the role of a missed parental figure or a romantic partner. The design of these bonds should not be treated as an afterthought or a side project, but as a central component of conversational interface design. This idea is not new in design: practitioners have long been advised not to design solely for “happy paths” (Wikipedia, 2025). However, the rise of LLMs has greatly expanded the variety and unpredictability of “unhappy paths.” These models are capable of interpreting diverse inputs and producing seemingly appropriate responses. As a result, designers must account for the fact that user inputs can emerge from many different needs.
Even with early chatbots before the era of LLMs, users often brought unexpected needs and behaviors. For example, people expressed love or tried to flirt with Siri, even though it was designed mainly to provide information and simple tasks (UNESCO and EQUALS Skills Coalition, 2019). With the arrival of LLMs, this baseline has grown more complex. LLMs are much stronger at understanding and generating natural text, so conversations feel more open and realistic. This creates new opportunities but also makes the challenge harder, since users can now say a much wider range of things—casual chatting, joking, flirting, threats, or even references to self-harm or harm to others. One recent case is Taco Bell’s drive-through chatbot, which was meant only to take orders. The system, however, had not been designed to handle cases where customers were impatient or playful. One customer ended up placing an order for 18,000 cups of water (McCallum, 2025).
This makes designers’ responsibilities heavier. They must not only build engaging and functional products but also encourage relationships that are constructive and supportive of user growth. Addressing these responsibilities requires iterative design processes, longitudinal studies of user experiences, and early involvement of stakeholders in the design process (Demiris, 2024; Durall Gazulla et al., 2023; Sadek et al., 2023). Not only should target users be consulted, but mental health professionals should also be part of the conversation from the outset (Yoo et al., 2025; Moylan and Doherty, 2025). Details such as tone of voice, word choice, style of expression, and even the persona and backstory created for a chatbot can carry deep meanings in the social relationships users form with it (Silva and Canedo, 2023; Chaves and Gerosa, 2021), and are therefore worthy of time and deliberate attention.
Adversarial attack-based security methods, such as Red Teaming (Shi et al., 2024; Nagireddy et al., 2024), can complement these design efforts. Many current approaches rely on other LLMs or automated identification systems (Shi et al., 2024), but these may miss the adversarial or playful cases teens might experiment with. To capture these behaviors, we recommend involving diverse groups of teens in testing and evaluation of chatbots. Designers of all conversational interfaces, regardless of their intended use, must integrate both sensitivity to user well-being and resilience against adversarial misuse.
5.2. Attachment-awareness: Design Without Assumptions
Our second principle encourages attachment-awareness and emphasizes how chatbot responses should be shaped according to those comprehensive needs stated in the first principle of our proposed framework. We propose this principle based on our finding that for some teens, chatbots acted as a coping mechanism. They turned to them out of loneliness, social isolation, or in search of mental health support. In some posts, users even sought comfort from a chatbot to fill the absence of a significant relationship in their lives, such as that between parent and child (§4.1.1). Attachment-aware design of such chatbots should be guided by the recognition that teens often turn to them to fill the void of relationships they lack in their lives. It is essential to attend to these needs rather than dismiss them.
Computer technologies are often shaped by the perspectives of people who are educated, financially secure, and accustomed to stable living conditions (Linxen et al., 2021). Some research assumed that most users grow up in stable households with parents who reliably meet their physical, psychological, and social needs, overlooking those whose experiences fall outside this norm (Marcu and Huh-Yoo, 2023). So, within chatbot design, this perspective can tempt designers to dismiss users’ attempts to form deeper connections with a default, impersonal message.
Attachment-informed design emphasizes that people’s lives span a spectrum of relationships—some supportive and others harmful (Marcu and Huh-Yoo, 2023). Designers should not assume that users already have healthy relationships in their lives, or that they can easily obtain them. In the context of chatbots, this means recognizing that not all teens grow up in families where parental love is reliable, and some may lack supportive connections with peers or access to appropriate mental health resources. It is understandable that such individuals may turn to chatbots to address unmet needs.
At the same time, many of these interactions resemble parasocial relationships, where users form one-sided bonds with media figures or, in this case, AI characters. For younger users, this possibility is even stronger. Developmental research shows that many children naturally engage in pretend play, create imaginary companions, or sustain friendships that exist primarily in their imagination (Calvert, 2017; Gleason and Hohmann, 2006). Teens, also, report that more than half have online-only friends, suggesting that digital or virtual relationships already feel natural to them (Lenhart et al., 2015). In this sense, forming a bond with a chatbot can be seen as a technological twist on pretend play (Severson and Woodard, 2018). Such experiences are not inherently harmful, but sometimes they blur the line between imagination, parasocial experience, and perceived social reality.
Designers must approach these situations with empathy, considering not only the immediate interaction but also the underlying needs that drive it. Simply rejecting users when they reach for more than a chatbot’s intended purpose risks worsening their sense of isolation. Our findings support this point: several teens expressed frustration when Character.AI introduced stricter filters, which made interactions feel impersonal and overly restricted (§4.3.2). This dismissive approach pushed users away from the platform but did not address the unmet needs that had led them there in the first place.
For children and adolescents, these parasocial or semi-social bonds may feel particularly real and emotionally deep (Cohen, 2004). This makes the task more complex: designers must account for the intensity of such relationships, and researchers must refine existing parasocial measures to better capture attachment, trust, conflict, and even antagonistic experiences with AI characters (Blake et al., 2025). Current scales and taxonomies, including parasocial theory, were not developed for these contexts and remain too limited (Blake et al., 2025). Parasocial theory assumes a lack of reciprocity, but this assumption is challenged by chatbots, especially those powered by generative AI. Even if one argues that reciprocity is absent because the non-human side has no awareness of the bond, the human user may still experience the relationship as genuine and perceive the chatbot as highly human-like. This concern is particularly acute for children and adolescents, who are still developing (Blake et al., 2025). Without a more nuanced understanding, we risk underestimating how strongly young users may invest in these relationships.
Research has shown that online safety technologies which rely heavily on external control—such as monitoring or content blocking—limit teens’ capacity for self-regulation or reflective decision-making (Wisniewski et al., 2017). A better approach is to address immediate concerns, such as preventing violence toward others or self-harm, and create a safe foundation from which users can build resilience and healthier coping strategies over time. Empathetic design and research, attentive to the realities of parasocial experience, can help ensure that chatbot interactions support rather than harm users’ development. How such empathetic interactions can be designed in practice is the focus of the next principle.
5.3. Respectful Empathy: More Than Kind Words
The third principle of our proposed framework calls for respectful empathy, which offers designers a foundation for creating chatbot responses that are both appropriate and supportive. We argue that chatbots should be designed to interact with users in ways that convey respect for their human dignity. Empathy in this sense does not mean abandoning honesty. Designers should not rely on superficial displays of empathy or anthropomorphic language that mislead users into thinking they are receiving genuine human care. Next, we explain how designers should draw on their own awareness and sensitivity to shape interactions that authentically communicate respect, as chatbots lack human understanding or emotions.
Our findings underline both the potential and the risk at stake. Many teens described chatbots as feeling so real that they became a central part of daily life (§4.2.2), improving mood and offering psychological comfort (§4.2.6). Yet this sense of realness was also fragile. Users were acutely aware that they were ultimately engaging with a machine, which created tension. Some wanted to disengage because the relationship felt artificial (§4.3.1), even though letting go was difficult (§4.2.5). Here lies both an opportunity and a threat. On the one hand, the conversational capabilities of chatbots can bring them close enough to users to affect well-being in meaningful ways. On the other hand, if this potential is used only superficially—by offering pleasant but hollow interactions—users may be drawn into relationships that feel comforting but are not grounded in genuine understanding or care. Respectful empathy means designers can bridge this gap, bringing their own insight and sensitivity into the design so that interactions are not merely scripted patterns of pleasant words but supportive exchanges that reinforce dignity and growth.
Attachment theory highlights two internal working models that shape mental health: how people see themselves and how they see others (Marcu and Huh-Yoo, 2023). In chatbot design, thoughtful use of language can strengthen both—helping users feel more capable and fostering a more positive view of relationships. Simply instructing a model to “sound empathetic” may produce sympathetic language, but it risks being superficial. If the goal is to support psychological well-being, designers should instead focus on reinforcing healthier internal models. For example, consistency is more powerful than flowery language. A prior study showed that users appreciated when chatbots remembered past conversations, such as recalling a health complaint mentioned earlier and following up on it (Jo et al., 2024). This kind of continuity gave users a stronger sense of being valued. Similarly, chatbots can use reflective statements that highlight users’ own progress. Messages that are more objective, such as, “You managed to keep working on your assignment even when it was difficult, which shows persistence,” encourage users to view themselves as capable without relying on anthropomorphic cues.
The final principle of attachment-informed design emphasizes engaging users without directly labeling interactions as mental health interventions. Across contexts, research indicated that barriers to seeking help for mental health include not recognizing the need for support, doubting that others can help, and fearing stigma (Wuthrich and Frei, 2015; Salaheddin and Mason, 2016; Aguirre Velasco et al., 2020). Designers who wish to create chatbots that genuinely care for users should take this into account. As we argued earlier, every interaction reflects unmet needs or relational gaps that users may be trying to fill through chatbots. Instead of focusing directly on mental health issues, designers can build experiences around activities, relationships, and communities. This is particularly important for teens, who often resist being told explicitly what to do. More effective interventions can guide them toward healthier ways of meeting their needs and filling relational gaps through social modeling, learning by example, and activities that they find intrinsically engaging (Marcu and Huh-Yoo, 2023).
5.4. Ease of Exit: Design for Disengagement
Our study showed that some teens recognized their own overuse of chatbots (§4.3.1), while others struggled to disengage despite wanting to (§4.2.5). This suggests that current systems do not sufficiently support user agency in disengagement and that design could provide better tools for users to manage and exit these relationships on their own terms. Designers must not only consider how interactions begin and how users can carry out their tasks effectively, but also how they can support an ease of exit. Generative AI poses particular challenges here: its ability to produce endless content on demand makes overuse especially likely. The final principle of our framework addresses this issue by proposing ways to support healthier disengagement.
What separates constructive from harmful disengagement is the degree of agency users retain (O’Brien et al., 2022). Current moderation strategies often rely on rigid restrictions, such as time limits, app locks, or usage reports (Team, 2024), and some jurisdictions have experimented with legal controls (Lomas, 2023). Yet such approaches can undermine autonomy and lead to negative experiences. In our study, some users reported abandoning platforms altogether after new censorship measures (§4.3.2), pointing to the unintended harms of abrupt, top-down interventions. Sudden restrictions may even cause users to feel that their AI companion has “died” or been replaced, leading to distress or grief (De Freitas et al., 2024).
A more effective path is to design disengagement mechanisms that preserve autonomy while taking advantage of the flexibility of language models. Prior research in video games, for example, showed that players felt less frustration when they were able to pause or save progress before leaving, rather than being forced to quit abruptly mid-quest (Tyack and Mekler, 2020; Davies and Blake, 2016). Translating this to chatbot design means that conversations should offer natural closure points, allowing users to complete a thought or goal before disengaging instead of being cut off suddenly and being able to save the state of the conversation.
Similarly, work on social media use among teens has shown that subtle nudges, delivered in casual and nonjudgmental language, can help users act more intentionally and reduce mindless engagement (Davis et al., 2023; Agha et al., 2025). Tools that help teens grow by managing risks, rather than shielding them through strict control, are more likely to be effective (Wisniewski et al., 2017). By contrast, systems grounded in surveillance or rigid restriction frequently erode trust and drive disengagement (Alelyani et al., 2019; Ghosh et al., 2018), especially since teens value privacy around emotionally sensitive topics (Dumaru and Al-Ameen, 2025). Building on these insights for chatbot design means avoiding blunt notifications or prescriptive commands. Instead, chatbots could weave supportive cues into everyday dialogue that encourage users to reconnect with offline life. During casual exchanges, the system might highlight progress the user has made, suggest opportunities for social contact, or propose light activities outside the platform. Interactions with chatbots could also be designed to support self-reflection by helping users name emotions, revisit meaningful events, and explore their thoughts in their own words, fostering emotional awareness and a stronger sense of self (Seo et al., 2024; Kim et al., 2024; Song et al., 2024).